rssLink RSS for all categories
 
icon_red
icon_green
icon_red
icon_red
icon_blue
icon_green
icon_green
icon_red
icon_red
icon_red
icon_orange
icon_green
icon_green
icon_green
icon_green
icon_blue
icon_green
icon_orange
icon_red
icon_green
icon_red
icon_red
icon_green
icon_red
icon_red
icon_red
icon_red
icon_orange
icon_green
 

FS#8703 — FS#12571 — Vps Cloud - SBG

Attached to Project— VPS
Incident
Infrastructure
CLOSED
100%
We are having problems of communication with a switch. We are intervening.
Date:  Monday, 09 February 2015, 10:02AM
Reason for closing:  Done
Comment by OVH - Saturday, 07 February 2015, 14:57PM

We restart the vcenter


Comment by OVH - Saturday, 07 February 2015, 14:58PM

Some VPS has still network problems. We are searching for the cause.


Comment by OVH - Monday, 09 February 2015, 09:51AM

Currently resynchronising the master and secondary switch


Comment by OVH - Monday, 09 February 2015, 09:52AM

There's no progress with the reconstruction. We will deploy a new switch.


Comment by OVH - Monday, 09 February 2015, 09:55AM

All the VMs are rebooting.
The deployment isn't fully complete. We will continue the reinstallation.


Comment by OVH - Monday, 09 February 2015, 09:56AM

Currently the robots are disabled on the cluster.


Comment by OVH - Monday, 09 February 2015, 09:58AM

We're still deploying a new switch. The operation is taking longer than anticipated.


Comment by OVH - Monday, 09 February 2015, 09:59AM

We've nearly completed the operation. We're currently verifying all the ACLs.


Comment by OVH - Monday, 09 February 2015, 10:02AM

The new switch is up. VMs have all been switched over. Everything is configured.